# Quantifying the Reliability of Predictions in Detection Transformers: Object-Level Calibration and Image-Level Uncertainty

### v2

@@ VI-A Proposed Method: Quantifying Reliability by Contrasting
VI-B Numerical Analysis
We compare different methods for quantifying per-image reliability based on different separation methods. For separation methods, we apply the optimal matching (as a reference) as well as the practical post-processing schemes such as fixed and adaptive confidence thresholding. We measure the Pearson correlation coefficient of each method with the $\mathsf{ImReli}$ computed based on AP metric (see Appendix for the detail).
@@ VI-B Numerical Analysis
Figure 6: Impact of the scaling factor ($\lambda$) on image-level UQ performance of $\mathsf{ContrastiveConf}$ (OCE). Pearson correlation coefficient (PCC) using various scaling factors is reported. The optimal scaling factor lies within the range of $5.0$ to $10.0$, while this range generalizes well across out-of-distribution datasets. Furthermore, it shows the efficacy of $\mathsf{ContrastiveConf}$ over $\mathsf{Conf}^{+}$ (i.e., $\mathsf{ContrastiveConf}$ with $\lambda=0.0$).
@@
In addition, we conduct an ablation study on the scaling factor ($\lambda$), with results presented in Figure 6. The study reveals that the best performing scaling factor for COCO dataset lies between $5.0$ and $10.0$. Notably, this range remains effective even for out-of-distribution datasets such as Cityscapes and Foggy Cityscapes. Furthermore, the results demonstrate that $\mathsf{ContrastiveConf}$ with $1.0\leq\lambda\leq 10.0$ consistently outperforms $\mathsf{Conf}^{+}$ (i.e., $\lambda=0.0$, the leftmost point on each line). However, we also observe a rapid drop in performance when the scaling factor is excessively large, particularly on the OOD dataset, Foggy Cityscapes. Thus, caution is needed when applying to substantially different datasets.
@@
Table 4: Comparison of post-processing schemes and baseline methods based on their Pearson correlation coefficients with image-level reliability when $\mathsf{Conf}^{+}$ and $\mathsf{ContrastiveConf}(\lambda=1.0)$ are applied, respectively. We evaluated three standard methods: confidence thresholding, top-$k$, and NMS (without confidence thresholding). For each scheme, the optimal hyperparameter selected from the validation set is shown in parentheses, and is applied on the test set. the strongest correlations are highlighted in bold. Confidence thresholding achieves the lowest OCE, demonstrating its better efficacy compared to the other schemes. For Deformable-DETR, Cal-DETR, and DINO, the optimal thresholds are approximately $0.3$, aligning well with values commonly employed in previous studies. For UP-DETR, the optimal thresholds often exceed $0.5$, highlighting the potential limitation of using a fixed threshold for image-level UQ.
@@
(a) Thresholding
(b) Top-$k$
(c) NMS
Figure 7: Impact of parameter selection on OCE (y-axis inverted) and the Pearson correlation coefficient (PCC) between $\mathsf{Conf}^{+}$ and image-level reliability in Cal-DETR on COCO for different post-processing schemes.
VII Comparative Study on Post-Processing
@@
Primarily, we evaluate OCE by varying the hyperparameter for each method (i.e., confidence threshold, $k$, and IoU threshold, respectively) and compare the resulting performance. In this first analysis, to exclude the dependency of the final performance on OCE, we choose the best hyperparameter on the validation set for each setting and compare their best possible performances. Table 3 shows that the confidence thresholding approach outperforms the top-$k$ and NMS approaches. The top-$k$ approach underperforms because the number of objects in an image varies significantly, making it impractical to determine a single optimal value for $k$. Top-$k$ becomes optimal when $k=20$. Given that the average and 95th percentile number of objects per image in the COCO dataset are 7 and 22, respectively, these results appear reasonable. Therefore, we confirm that using an excessively large number (e.g., 100) for top-$k$ is inadequate for achieving well-calibrated predictions. On the other hand, with NMS (when applied without preceding confidence thresholding), we empirically observe that although most of the optimal positive queries are included, a substantial number of negative queries are retained, often outnumbering the positive ones by several times.
@@
Lastly, we plot the OCE and image-level UQ performance with different hyperparameter selections to show the sensitivity of each scheme to the choice of hyperparameter. Figure 7 highlight that carefully identifying a reliable subset is crucial for achieving both high object-level calibration quality and effective image-level uncertainty quantification performance. We also present this with exemplary visualizations in Figure 8 and in the Appendix.
(a) Thresholding (0.1)
(b) Thresholding (0.3)
(c) Thresholding (0.9)
(d) Top-$1$
(e) Top-$20$
(f) Top-$100$
(g) NMS ($\tau=0.1$)
(h) NMS ($\tau=0.5$)
(i) NMS ($\tau=0.9$)
Figure 8: Exemplary visualization demonstrating the impact of parameter selection on the final subset of predictions in Cal-DETR for different post-processing schemes. Optimal positive and negative predictions are highlighted with green and red boxes, respectively. As shown, the top-$k$ and NMS approaches often include too many negative predictions, degrading the calibration quality. Confidence thresholding with too low of a threshold faces a similar issue, while too high of a threshold risks omitting positive predictions. Therefore, accurately identifying a reliable set of predictions significantly affects the reliability of DETR for downstream applications.
VIII Conclusion
The main contribution of our work is an in-depth analysis of the reliability of DETR frameworks. In particular, we reveal that DETR’s predictions for a given image exhibit varying calibration quality, highlighting the importance of identifying well-calibrated positive predictions. To address this challenge, we introduced a systematic framework that leverages our object-level calibration error metric to discern these positive predictions effectively. Furthermore, we proposed a novel uncertainty quantification method for estimating image-level reliability in DETR and conducted comparative studies of various post-processing schemes regarding their impact on DETR’s reliability. We hope our efforts expand the scope of DETR applications by enabling more precise and reliable deployment, sparking further research into the area of positive prediction identification.
Acknowledgments
Appendix A Dataset
Appendix B Performance Metrics
Average precision (AP) [17] is the most common accuracy metric for object detectors. Before calculating AP, the top-$k$ (e.g., typically $100$) class predictions for a given image are kept and sorted, while the others are discarded. Then, in order of this sorting, the predictions are compared against against the ground-truth bounding boxes, and the predicted label is compared with the ground truth label if the IoU is large enough. The precision/recall curve is calculated over the sorted predictions, and AP averages the precision across a set of 101 evenly spaced recall thresholds:
$\text{AP}=\frac{1}{101}\sum_{r\in[0:0.01:1]}p(r)$ (16)
where $p(r)$ is calculated by the maximum precision at recall level $r$.
[…]
$\text{LRP}_{\tau}=\frac{1}{N_{\text{FP}}+N_{\text{FN}}+N_{\text{TP}}}\left(N_{\text{FP}}+N_{\text{FN}}+\sum_{i=1}^{N_{\text{TP}}}\frac{1-\text{IoU}_{i}}{1-\tau}\right)$ (17)
where $\tau$ is IoU threshold, $i$ denotes the index of each true positive prediction, and IoUi represents the IoU between that prediction and the best-matching ground-truth object. LRP is then computed by averaging LRP0.5 and LRP0.75.
Appendix C Omitted Experimental Results
This section provides the omitted experimental results. Figure 10 extends Figure 5 in Section V-A across different DETR variants and datasets. Figure 11 extends Figure 8 in Section VII across different DETR variants and post-processing schemes.
Appendix D Contrasting with Confounding Negatives for Image-level UQ
In Table 2, we observe an interesting empirical result: $\mathsf{Conf}^{+}$ using the top 100 predictions (i.e., the fourth row) exhibits a stronger negative correlation compared to $\mathsf{Conf}^{-}$ using the optimal negative predictions (i.e., the second row). While a stronger negative correlation might not be desirable when used alone for UQ, this phenomenon can be leveraged within our contrasting framework, similar to how negative predictions are utilized.
For instance, rather than contrasting the average confidence scores of positive predictions against those of negative (non-positive) predictions, we can instead contrast them against the top 100 predictions. However, since positive predictions are included within this top 100 set, an alternative and potentially clearer approach is to contrast the positive predictions specifically against the top 100 non-positive predictions.
Nevertheless, employing a fixed value of $k=100$ might introduce unintended consequences, particularly depending on the number of ground truth objects present in an image. We hypothesize that among non-positive predictions, those with bounding boxes significantly overlapping the positive predictions play a more meaningful role. Consequently, negative predictions that have random bounding boxes and very low confidence scores (as illustrated in Figure 1a) can be safely disregarded during the contrastive evaluation.
To formalize this intuition, we introduce the concept of confounding negatives:
Definition 3.
Confounding negative predictions are defined as non-positive predictions whose bounding boxes significantly overlap with positive predictions, specifically exceeding a threshold overlap $\delta$:
$\hat{\mathcal{D}}_{\theta}^{-\delta}(\bm{x})=\Big{\{}d^{-}\in\hat{\mathcal{D}}_{\theta}^{-}(\bm{x})\mid\max_{d^{+}\in\hat{\mathcal{D}}_{\theta}^{+}(\bm{x})}\mathrm{IoU}(d^{-},d^{+})\geq\delta\Big{\}}.$ (18)
We conduct an ablation study comparing the original approaches, which use the entire set of negative predictions, against variants that employ different subsets of negatives described above. Our findings are summarized in Table 5.
Overall, we see that restricting attention to negatives with significant overlap can sharpen the negative correlation for $\mathsf{Conf}^{-}$, especially on in-distribution data (COCO), supporting our hypothesis that misleading or “confouding” boxes are more informative for contrastive scoring. On the other hand, using entire negatives sometimes proves more robust in near-OOD and OOD scenarios, particularly for $\mathsf{ContrastiveConf}$.
Table 5: Comparison of correlation results under $\mathsf{Conf}^{-}$ (rows 1–5) and $\mathsf{ContrastiveConf}$ (rows 6–10) when using different negative sets on three datasets: COCO (in-distribution), Cityscapes (near OOD), and Foggy Cityscapes (OOD). “Entire Negatives” refers to all non-positive boxes, while “Top-100 Negatives” and “Top-100 Positives” limit the selection to the 100 highest-confidence predictions in each category. Confounding negatives restricts the negative set to boxes overlapping positive boxes by at least $\delta$. Stronger (more negative) correlation values under $\mathsf{Conf}^{-}$ appear when negatives more closely resemble the positive boxes, and $\mathsf{ContrastiveConf}$ tends to benefit from using such confounding negatives, especially in in-distribution scenarios.
Methods Negatives COCO (in-distribution) Cityscapes (near OOD) Foggy Cityscapes (OOD)
$\mathsf{Conf}^{-}$ Entire Negatives -0.640 -0.575 -0.571 -0.585 -0.349 -0.382 -0.379 -0.394 -0.256 -0.322 -0.253 -0.283
Top-100 Positives -0.619 -0.603 -0.581 -0.601 -0.409 -0.374 -0.372 -0.391 -0.262 -0.270 -0.239 -0.229
Top-100 Negatives -0.640 -0.607 -0.607 -0.620 -0.349 -0.385 -0.415 -0.404 -0.256 -0.315 -0.272 -0.279
Confounding Negatives ($\delta=0.5$) -0.641 -0.640 -0.574 -0.647 -0.330 -0.351 -0.364 -0.386 -0.250 -0.260 -0.254 -0.261
Confounding Negatives ($\delta=0.75$) -0.635 -0.628 -0.561 -0.602 -0.385 -0.388 -0.384 -0.390 -0.285 -0.279 -0.269 -0.254
$\mathsf{ContrastiveConf}$ Entire Negatives 0.656 0.588 0.628 0.613 0.359 0.407 0.441 0.440 0.275 0.350 0.317 0.327
Top-100 Positives 0.632 0.610 0.604 0.619 0.412 0.379 0.388 0.404 0.270 0.277 0.257 0.247
Top-100 Negatives 0.656 0.612 0.629 0.633 0.359 0.401 0.439 0.431 0.275 0.332 0.299 0.308
Confounding Negatives ($\delta=0.5$) 0.653 0.641 0.623 0.654 0.340 0.371 0.434 0.418 0.268 0.284 0.314 0.296
Confounding Negatives ($\delta=0.75$) 0.647 0.630 0.613 0.619 0.388 0.406 0.448 0.418 0.298 0.301 0.326 0.286
(a) Object-level Calibration Error
(b) Object-level Calibration Error
(Foggy Cityscapes)
(c) Corr. to Image-level Reliability
(d) Corr. to Image-level Reliability
(Foggy Cityscapes)
Figure 9: A visualization of the difference in calibration between positive and negative predictions on the Cityscapes and Foggy Cityscapes datasets. In Figures 9a and 9b, Object-Level Calibration Error (OCE) values of their confidence scores are shown, where a lower score represents better-calibrated predictions. In Figures 9c and 9d, Pearson Correlation Coefficient (PCC) values between the ground truth $\mathsf{ImReli}$ and the average confidence scores are shown, where a higher value is better.

### v3

@@ VI-B Proposed Method: Quantifying Reliability by Contrasting
VI-C Numerical Analysis
The model is tested on three datasets with varying levels of out-of-distribution (OOD) characteristics: COCO (in-distribution), Cityscapes (near OOD), and Foggy Cityscapes (OOD).
We compare different methods for quantifying per-image reliability based on different separation methods. For separation methods, we apply the optimal matching (as a reference) as well as the practical post-processing schemes such as fixed and adaptive confidence thresholding. We measure the Pearson correlation coefficient of each method with the $\mathsf{ImReli}$ computed using the standard AP metric.
@@ VI-C Numerical Analysis
Figure 12: Impact of the scaling factor ($\lambda$) on image-level UQ performance of $\mathsf{ContrastiveConf}$ (OCE). Pearson correlation coefficient (PCC) using various scaling factors is reported on (a) In-distribution (COCO), (b) Near OOD (Cityscapes), and (c) Far OOD (Foggy Cityscapes). The optimal scaling factor lies within the range of $5.0$ to $10.0$, while this range generalizes well across out-of-distribution datasets. Furthermore, it shows the efficacy of $\mathsf{ContrastiveConf}$ over $\mathsf{Conf}^{+}$ (i.e., $\mathsf{ContrastiveConf}$ with $\lambda=0.0$).
@@
In addition, we conduct an ablation study on the scaling factor ($\lambda$), with results presented in Figure 12. The study reveals that the best performing scaling factor for COCO dataset lies between $5.0$ and $10.0$. Notably, this range remains effective even for out-of-distribution datasets such as Cityscapes and Foggy Cityscapes. Furthermore, the results demonstrate that $\mathsf{ContrastiveConf}$ with $1.0\leq\lambda\leq 10.0$ consistently outperforms $\mathsf{Conf}^{+}$ (i.e., $\lambda=0.0$, the leftmost point on each line). That said, we also observe a rapid drop in performance when the scaling factor is excessively large, particularly on the OOD dataset, Foggy Cityscapes. Thus, we do not recommend using large scaling factors or relying solely on the negative predictions.
VII Conclusion
The main contribution of our work is an in-depth analysis of the reliability of DETR frameworks. In particular, we reveal that DETR’s predictions for a given image exhibit varying calibration quality, highlighting the importance of identifying well-calibrated predictions. To address this challenge, we introduced a systematic framework that leverages our object-level calibration error metric to discern these reliable predictions effectively. Furthermore, we proposed a novel uncertainty quantification method for estimating image-level reliability in DETR and conducted comparative studies of various post-processing schemes regarding their impact on DETR’s reliability. We hope our efforts expand the scope of DETR applications by enabling more precise and reliable deployment, sparking further research into the area of reliable positive prediction identification.
Acknowledgments
Appendix A Dataset
Appendix B Detailed Theoretical Proofs
B-A Proof of Proposition 1
Proposition 1.
The minimum achievable expected total loss is lower bounded by the entropy of the ground-truth class distribution, i.e.,
$\mathbb{E}\left[\mathcal{L}_{\text{total}}\right]\geq\mathcal{H}(p_{\text{gt}})+\min_{\hat{b}}\mathbb{E}[\mathcal{L}_{\text{box}}(\hat{b},b_{\text{gt}})],$
where $\mathcal{H}(p_{\text{gt}})=-\sum_{c\in\mathcal{C}}p_{\text{gt}}(c)\log(p_{\text{gt}}(c))$.
Proof.
The expected total loss is the sum of the expected loss for the optimal positive query matched to $y_{\text{gt}}$ (denoted as index $p$) and the expected losses for the M1M-1 remaining queries matched to $\emptyset$ (denoted as indices $\{n\neq p\}$):
$\mathbb{E}[\mathcal{L}_{\text{total}}]=\mathbb{E}[\mathcal{L}(\hat{y}_{p},y_{\text{gt}})]+\sum_{n\neq p}\mathcal{L}(\hat{y}_{n},\emptyset).$
From Equation 4, $\mathcal{L}(\hat{y}_{n},\emptyset)=-w_{\emptyset}\log(\hat{p}_{n}(\emptyset))$; and since $\log(\hat{p}_{n}(\emptyset))\leq 0$: $\mathcal{L}(\hat{y}_{n},\emptyset)\geq 0$. Therefore, the total expected loss is bounded below by:
$\mathbb{E}[\mathcal{L}_{\text{total}}]\geq\mathbb{E}[\mathcal{L}(\hat{y}_{p},y_{\text{gt}})].$
From Equation IV-C, we have:
$\mathbb{E}[\mathcal{L}(\hat{y}_{p},y_{\text{gt}})]=\underbrace{-\sum_{c\in\mathcal{C}}p_{\text{gt}}(c)\log(\hat{p}_{p}(c))}_{\mathbb{E}[\mathcal{L}_{\text{cls}}]}+\mathbb{E}_{b_{\text{gt}}\sim\mathcal{B}}[\mathcal{L}_{\text{box}}(\hat{b}_{p},b_{\text{gt}})].$
We minimize the two terms on the right-hand side independently. First, we find the minimum classification loss $\mathbb{E}[\mathcal{L}_{\text{cls}}]$. We define $S=\sum_{c\in\mathcal{C}}\hat{p}_{p}(c)=1-\hat{p}_{p}(\emptyset)$ and $\hat{q}_{p}(c)=\hat{p}_{p}(c)/S$. Then:
$\mathbb{E}[\mathcal{L}_{\text{cls}}]$ $=-\sum_{c\in\mathcal{C}}p_{\text{gt}}(c)(\log(S)+\log(\hat{q}_{p}(c)))$
$=-\log(S)+\mathrm{CE}(p_{\text{gt}},\hat{q}_{p}),$
where $\mathrm{CE}(p_{\text{gt}},\hat{q}_{p})=\sum_{c\in\mathcal{C}}p_{\text{gt}}(c)\log(\hat{q}_{p}(c))$ is the cross-entropy. By Gibbs’ inequality, $\mathrm{CE}(p_{\text{gt}},\hat{q}_{p})\geq\mathcal{H}(p_{\text{gt}})$ and equality holds if and only if $\hat{q}_{p}=p_{\text{gt}}$. Additionally, $\log(S)\geq 0$ and equality holds if and only if $S=1$ (i.e., $\hat{p}_{p}(\emptyset)=0$).
Denoting the minimum box loss as $\min_{\hat{b}}\mathbb{E}[\mathcal{L}_{\text{box}}(\hat{b},b_{\text{gt}})]$, a lower bound of $\mathbb{E}_{b_{\text{gt}}\sim\mathcal{B}}[\mathcal{L}_{\text{box}}(\hat{b}_{p},b_{\text{gt}})]$, we have:
$\mathbb{E}[\mathcal{L}_{\text{total}}]\geq\mathcal{H}(p_{\text{gt}})+\min_{\hat{b}}\mathbb{E}[\mathcal{L}_{\text{box}}(\hat{b},b_{\text{gt}})].$
B-B Proof of Proposition 2
Proposition 2 (Optimal Prediction Set).
The unique globally optimal prediction set that minimizes $\mathbb{E}\left[\mathcal{L}_{\text{total}}\right]$ is $\{(\hat{p}^{*}_{i},\hat{b}^{*}_{i})\}_{i=1}^{M}$, where
  1. 1.
    for the optimal positive prediction ($i=p$): $\hat{p}^{*}_{i}=p_{\text{gt}}$ (i.e., $\hat{p}^{*}_{i}(c)=p_{\text{gt}}(c)$ $\forall c\in\mathcal{C}$ and $\hat{p}^{*}_{i}(\emptyset)=0$) and the box $\hat{b}^{*}_{i}$ minimizes $\mathbb{E}[\mathcal{L}_{\text{box}}(\hat{b}_{i},b_{\text{gt}})]$, and
  2. 2.
    for any optimal negative prediction ($i\neq p$): $\hat{p}^{*}_{i}(c)=0$ $\forall c\in\mathcal{C}$ and $\hat{p}^{*}_{i}(\emptyset)=1$.
Proof.
We evaluate the expected total loss for the proposed strategy $\{(\hat{p}^{*}_{i},\hat{b}^{*}_{i})\}_{i=1}^{M}$:
  1. 1.
    For the optimal positive prediction: The strategy sets $\hat{p}^{*}_{p}=p_{\text{gt}}$ and $\hat{b}^{*}_{p}$ to minimize the expected box loss. The resulting expected loss is:
    $\mathbb{E}[\mathcal{L}(\hat{y}^{*}_{p},y_{\text{gt}})]$ $=\min_{\hat{b}}\mathbb{E}[\mathcal{L}_{\text{box}}(\hat{b},b_{\text{gt}})]-\sum_{c\in\mathcal{C}}p_{\text{gt}}(c)\log(p_{\text{gt}}(c))$
    $=\min_{\hat{b}}\mathbb{E}[\mathcal{L}_{\text{box}}(\hat{b},b_{\text{gt}})]+\mathcal{H}(p_{\text{gt}}).$
  2. 2.
    For any optimal negative prediction: The strategy sets $\hat{p}^{*}_{n}(\emptyset)=1$. The resulting expected loss is:
    $\mathcal{L}(\hat{y}^{*}_{n},\emptyset)=-w_{\emptyset}\log(1)=0.$
Summing these terms, the strategy achieves the lower bound derived in Proposition 1:
$\mathbb{E}[\mathcal{L}_{\text{total}}]=\min_{\hat{b}}\mathbb{E}[\mathcal{L}_{\text{box}}(\hat{b},b_{\text{gt}})]+\mathcal{H}(p_{\text{gt}}).$
Any deviation from this strategy results in a strictly higher loss:
  • If $\hat{p}_{n}(\emptyset)<1$, then $w_{\emptyset}\log(\hat{p}_{n}(\emptyset))>0$.
  • If $\hat{b}_{p}$ is not the minimizer, the box loss term increases.
  • If $\hat{p}_{p}\neq p_{\text{gt}}$, then either $\log(S)>0$ or $\mathrm{CE}(p_{\text{gt}},\hat{q}_{p})>\mathcal{H}(p_{\text{gt}})$ (by Gibbs’ inequality).
Thus, the proposed prediction set is the unique global minimizer. ∎
B-C Proof of Proposition 3
Proposition 3 (Optimal Strategy under Uncertainty).
For a query with a matching probability $P_{i}$ to the ground-truth object whose class distribution is $p_{\text{gt}}$, the class prediction that minimizes the expected loss is $\hat{p}_{i}^{*}$, where
  1. 1.
    the prediction for any foreground class $c\in\mathcal{C}$ is:
    $\hat{p}_{i}^{*}(c)=p_{\text{fg}}^{*}(P_{i})p_{\text{gt}}(c).$ (9)
  2. 2.
    the prediction for the background class is
    $\hat{p}_{i}^{*}(\emptyset)=1-p_{\text{fg}}^{*}(P_{i}),$
where $p_{\text{fg}}^{*}(P_{i})$ is the optimal foreground probability, defined as
$p_{\text{fg}}^{*}(P_{i})=\frac{P_{i}}{P_{i}+w_{\emptyset}(1-P_{i})}.$ (10)
Proof.
We find the optimal probability distribution $\hat{p}_{i}$ by minimizing the expected classification objective:
$J(\hat{p}_{i})=-P_{i}\sum_{c\in\mathcal{C}}p_{\text{gt}}(c)\log(\hat{p}_{i}(c))-w_{\emptyset}(1-P_{i})\log(\hat{p}_{i}(\emptyset)).$
This is a constrained optimization problem, as we must satisfy the constraint that:
$g(\hat{p}_{i})=\sum_{c\in\mathcal{C}}\hat{p}_{i}(c)+\hat{p}_{i}(\emptyset)-1=0.$
We form the Lagrangian $\mathcal{J}$ with a Lagrange multiplier $\lambda$:
$\mathcal{J}=J(\hat{p}_{i})-\lambda g(\hat{p}_{i}).$
We find the optimum by setting the partial derivatives of $\mathcal{J}$ with respect to each probability $\hat{p}_{i}(c)$ to zero.
For any foreground class $c\in\mathcal{C}$:
$\frac{\partial\mathcal{J}}{\partial\hat{p}_{i}(c)}=0\implies\hat{p}_{i}(c)=-\frac{P_{i}\cdot p_{\text{gt}}(c)}{\lambda}.$
For the background class $\emptyset$:
$\frac{\partial\mathcal{J}}{\partial\hat{p}_{i}(\emptyset)}=0\implies\hat{p}_{i}(\emptyset)=-\frac{w_{\emptyset}(1-P_{i})}{\lambda}.$
We solve for $\lambda$ by substituting these back into the constraint $g(\hat{p}_{i})=0$:
$\sum_{c\in\mathcal{C}}\left(-\frac{P_{i}\cdot p_{\text{gt}}(c)}{\lambda}\right)+\left(-\frac{w_{\emptyset}(1-P_{i})}{\lambda}\right)=1.$
Since $\sum_{c\in\mathcal{C}}p_{\text{gt}}(c)=1$ by definition:
$\frac{P_{i}}{\lambda}-\frac{w_{\emptyset}(1-P_{i})}{\lambda}=1\implies\lambda=-(P_{i}+w_{\emptyset}(1-P_{i})).$
Finally, we substitute this $\lambda$ back to find the optimal predictions $\hat{p}_{i}^{*}$:
$\hat{p}_{i}^{*}(c)=-\frac{P_{i}\cdot p_{\text{gt}}(c)}{-(P_{i}+w_{\emptyset}(1-P_{i}))}=\left(\frac{P_{i}}{P_{i}+w_{\emptyset}(1-P_{i})}\right)p_{\text{gt}}(c),$
and
$\hat{p}_{i}^{*}(\emptyset)=-\frac{w_{\emptyset}(1-P_{i})}{-(P_{i}+w_{\emptyset}(1-P_{i}))}=1-\left(\frac{P_{i}}{P_{i}+w_{\emptyset}(1-P_{i})}\right).$
Appendix C Omitted Experimental Results
This section provides the omitted experimental results. Figure 14 extends Figure 8 across different DETR variants and datasets. Figure 13 extends Figure 5 across different DETR variants and datasets.
Figure 13: Empirical cumulative distribution function (ECDF) plots for maximum confidence score and maximum IoU, comparing optimal positive predictions (blue) and optimal negative predictions (red) on the COCO dataset.
@@ Appendix C Omitted Experimental Results
Table 4: Comparison of post-processing schemes and baseline methods based on their Pearson correlation coefficients with image-level reliability when $\mathsf{Conf}^{+}$ and $\mathsf{ContrastiveConf}(\lambda=1.0)$ are applied, respectively. We evaluated three standard methods: confidence thresholding, top-$k$, and NMS (without confidence thresholding). For each scheme, the optimal hyperparameter selected from the validation set is shown in parentheses and is applied on the test set. The strongest correlations are highlighted in bold. Confidence thresholding achieves the lowest OCE, demonstrating its better efficacy compared to the other schemes. For Deformable-DETR, Cal-DETR, and DINO, the optimal thresholds are approximately $0.3$, aligning well with values commonly employed in previous studies. For UP-DETR, the optimal thresholds often exceed $0.5$, highlighting the potential limitation of using a fixed threshold for image-level UQ.
@@
Appendix D Comparative Study on Post-Processing
@@
Primarily, we evaluate OCE by varying the hyperparameter for each method (i.e., confidence threshold, $k$, and IoU threshold, respectively) and compare the resulting performance. In this first analysis, to exclude the dependency of the final performance on OCE, we choose the best hyperparameter on the validation set for each setting and compare their best possible performances. Table 3 shows that the confidence thresholding approach outperforms the top-$k$ and NMS approaches. The top-$k$ approach underperforms because the number of objects in an image varies significantly, making it impractical to determine a single optimal value for $k$. Top-$k$ becomes optimal when $k=20$. Given that the average and 95th percentile number of objects per image in the COCO dataset are 7 and 22, respectively, these results appear reasonable. Therefore, we confirm that using an excessively large number (e.g., 100) for top-$k$ is inadequate for achieving well-calibrated predictions. On the other hand, with NMS (when applied without preceding confidence thresholding), we empirically observe that although most of the optimal positive predictions are included, a substantial number of negative predictions are retained, often outnumbering the positive ones by several times.
@@
Lastly, we plot the OCE and image-level UQ performance with different hyperparameter selections to show the sensitivity of each scheme to the choice of hyperparameter. Figure 15 highlight that carefully identifying a reliable subset is crucial for achieving both high object-level calibration quality and effective image-level uncertainty quantification performance. We also present this with exemplary visualizations in Figure 16.
@@
(a) Thresholding (0.1)
(b) Thresholding (0.27)
(c) Thresholding (0.9)
(d) Top-11
(e) Top-2020
(f) Top-$100$
(g) NMS ($\tau=0.1$)
(h) NMS ($\tau=0.5$)
(i) NMS ($\tau=0.9$)
Figure 16: Exemplary visualization demonstrating the impact of parameter selection on the final subset of predictions in Cal-DETR for different post-processing schemes. Optimal positive and negative predictions are highlighted with green and red boxes, respectively. As shown, the top-$k$ and NMS approaches could include too many negative predictions, degrading the calibration quality. Confidence thresholding with too low of a threshold faces a similar issue, while too high of a threshold risks omitting positive predictions. Therefore, accurately identifying a reliable set of predictions significantly affects the reliability of DETR for downstream applications.